A buffered solution containing dissolved aniline, C6H5NH2, and aniline hydrochloride, C6H5NH3Cl, has a pH of 5.65.

a) Determine the concentration of C6H5NH3 in the solution if the concentration of C6H5NH2 is 0.230 M. The pKb of aniline is 9.13.

[C6H5NH3+] = ______M

b)Calculate the change in pH of the solution, ?pH, if 0.386 g NaOH is added to the buffer for a final volume of 1.45 L. Assume that any contribution of NaOH to the volume is negligible.

\DeltapH = ________
Answer #1

a) it represent basic buffer

pOH=pKb+log(C6H5NH3/C6H5NH2)

14-5.65=9.13+log(C6H5NH3/C6H5NH2)

log(C6H5NH3/C6H5NH2)=-0.78

(C6H5NH3/C6H5NH2)=0.166

[C6H5NH3]=0.166×0.230=0.0382M

b) [NaOH] added=0.386/40×1.45=0.00536M

It would reduce concentration of C6H5NH3 and increase concentration of C6H5NH2.

Now [C6H5NH3]=0.0382-0.00536=0.0328

[C6H5NH2]=0.230+0.0536=0.23536

pOH=pKb+log(C6H5NH3/C6H5NH2)=9.13+log(0.0328/0.23536)=8.274

pH=5.726

Change in pH=0.0726